### v1

@@ 1 Introduction
Employing Artificial Intelligence (AI) for automation of manufacturing has resulted in increased efficiency, precision, and flexibility and created a pardigm shift in the design of manufacturing systems. AI has been successfully applied to a vast array of manufacturing tasks in the industry [1, 2]. While AI-based methods have improved the manufacturing process, there are still challenges in ensuring that the AI-based black box systems continue to be reliable and robust. The most fundamental layer underlying all smart manufacturing systems is object detection, which allows the system to identify the type and position of the objects that it needs to handle. Object detection is a established computer vision problem, which involves identifying and categorizing specific objects of interest within a larger image by placing a bounding box around each detected object [3].
An effective automation system also requires proper sensor design to provide adequate coverage over the environment and allow the system to properly observe the scene and the objects [4]. The past decades have seen great advancement in sensor hardware. Camera resolutions have increased and they have become more affordable at the same time. Similarly, consumer applications have facilitated mass manufacturing of 3D sensors like lidars and stereo cameras, which are great sensors for smart manufacturing. Despite these improvements, sensors have inherent limitations rooted in their physics. For example, a single image captured by camera does not carry depth information. 3D sensors capture point cloud data which addresses this issue, but these sensors are typically low resolution and do not provide color information.
An effective and reliable automation system requires selecting the right sensors and an object detection system that can effectively ingest the data provided by these sensors. Prior work has shown the limitations of object detection systems that rely solely on cameras [5, 6, 7, 8]. Image distortions like blur and noise can significantly lower the detection accuracy [5]. Although cameras provide color information, there are environments where there is low contrast between objects and the background, and as a result detection accuracy may suffer [9] and cameras may not be enough for handling these environments. Moreover, camera-based object detection systems are sensitive to illumination and can be fragile if there are changes to the lighting conditions. It has also been shown that illumination can negatively affect [6] camera-based object detection systems, because they struggle to generalize to operate under lighting conditions different from what they have experienced during training. Similarly, the performance of these systems diminishes in scenarios where objects vary significantly in size or when they blend indistinguishable with the background in camera’s view [7].
@@
However, none of the existing methods have explored efficient four-channel RGB+D inputs in the context of object detection tasks. The processing of RGB and depth data through a single shared backbone offers several advantages. Firstly, when RGB and depth data are stacked at the input level, the model is able to efficiently extract features that merge information from both modalities, for example color and texture from RGB alongside spatial and structural details from depth. This integrated approach enables the network to learn more meaningful features early in the feature extraction process. By processing both data types simultaneously, the network can better generalize to unseen data as it learns to recognize and interpret patterns across both modalities. Additionally, employing a single backbone for processing reduces the computational overhead significantly. This computational efficiency comes from managing only one set of weights during back-propagation and inference. This approach creates a leaner model which is faster and less costly to run in a manufacturing plant. Lastly, the lower architectural complexity and lack of complex fusion mechanisms simplifies the development and tuning process for training and deploying the final model in the manufacturing environment.
This work introduces a novel method that employs early sensor fusion, where data from RGB images and 3D point clouds are integrated at the data level before being used by the detection algorithm. This method utilizes the combined strengths of both modalities for a more comprehensive representation of the scene that enhances object detection capabilities. Early fusion facilitates the extraction of comprehensive features that embody both the visual and spatial attributes of objects, which addresses the challenges posed by occlusions, variable sizes, and complex background information more effectively than late fusion or single-sensor methods. Hence, this work presents a novel model tailored for processing RGB-D images for object detection tasks in manufacturing. The rest of the paper is organized as follows. Section  2 discusses the methodology of the multimodal object detection framework. Section  3 outlines experiments which compare this approach to both RBG-only and Depth-only object detection of a NIST manufacturing task board. Section  4 discusses the experiment results, and Section  5 discusses the conclusions and suggestions for future work.
@@ 3.1 Assembly Task Board
The task board used in this work is derived from the modified NIST task board configuration which contains nine components serving as the training and testing ground for the proposed object detection system. This board which is shown in Figure 2 includes a large gear, a small gear, a USB-C connector, a nut, a waterproof connector, pairs of small and large rectangular pins, and pairs of small and large round pins. The usage of this task board introduces realistic scenarios where lighting conditions, component orientation, and material finishes can affect the accuracy of object identification. This selection of components provides a comprehensive framework for evaluating the capabilities of the system under study. Variations in surface texture, from gear smoothness to the metallic finish of connectors, generate reflections and shadows that may obscure features or mimic other objects, risking misclassification. In addition, material properties including pin translucency and intricate details of components such as USB-C connectors introduce visual variability that demands sophisticated interpretation by the network. These factors underscore the need for an advanced, adaptable neural network and thorough pre-processing to ensure accurate detection amidst the multifaceted visual conditions of the task board.
@@ 3.2 Sensors and Calibration
Рис. 3: Multi-sensor setup featuring a Basler RGB camera and an Intel RealSense camera mounted to the robot’s table.
The object detection framework ingests data from two independent sensors capturing image and depth data. Figure 3 shows the multi-sensor setup used in our experiments. A Basler camera captures images, while an Intel RealSense depth camera senses the world in 3D in the form of point clouds. The sensors are mounted together and affixed to the robot’s table. This combination of sensors was selected to provide visual details of the scene in addition to the spatial relationships and dimensions of the objects under study. Although certain types of hardware exist which can provide both RGB and depth data from a single sensor, our multi-sensor framework provides greater compatibility and generalizability to diverse manufacturing problems and environments.
The two sensors need to be calibrated against each other since the assembly task and the object detection model require precise alignment between the image and depth data received from the environment. Camera calibration [23] involves the precise estimation of camera parameters, including intrinsic and extrinsic, to infer accurate geometric features from captured sequences. To conduct the calibration, an asymmetric 10x7 checkerboard pattern was placed in various positions and orientations within the two sensor’s field of view and 60 pairs of image and point clouds were captured. MATLAB’s Lidar-Camera Calibration application [24] was used to tune the calibration parameters, as shown in Fig. LABEL:fig:calibration. Fig. LABEL:fig:calibration shows that the camera intrinsics are properly calibrated, therefore the red boundary perfectly lines up with the checkerboard’s perimeter. However, the extrinsics are not calibrated yet, so the point cloud does not align with the checkerboard in the camera view. At the end of the process, the visual display and the error metrics signal a proper calibration. Hence, this process involved iteratively removing image and point cloud pairs with high calibration errors until the maximum translation error was below $0.0045m$ and the maximum rotation error was below $4.5\degree$. At this point, a total of 37 image - point cloud pairs were used to compute the calibration parameters.
@@ 3.3 Generating Depth Maps
Since the model receives RGB images encoded as unsigned integers in range $[0,255]$, this work further scale the depth values as
@@
where $D_{s}$ is the scaled depth value. After this transformation, the depth map is concatenated with the RGB image to form a four-channel image where all values range consistently from 0 to 255, which makes the data suitable for feeding to the model. The four-channel RGB-D images can be conveniently precomputed and stored to disk since many image formats like PNG support a fourth alpha channel mainly used for encoding transparency information for each pixel.
@@
Рис. 4: The three variants of the object detection model used in experiments. a. RGB-only: The model receives only camera images and predicts object class scores and bounding boxes for the detected objects. b. RGBD-Man: The main model which concatenates RGB and depth information into a four-channel input and predicts objects from the fused inputs. c. Depth-only: The model which predicts objects given only a single-channel depth map of scene.
@@ 3.4 Data Collection and Labeling
A dataset containing 301 pairs of images and associated point clouds captured from varied configurations of the assembly components on and off the assembly board were created. Placing the objects on and off the board enables the robot to detect and localize the components both before and after they get installed on the assembly board. The position and orientation of the assembly board within the sensors’ field of view were also varied so that the model can generalize to varied task configurations. It was also ensured that each class of components is adequately present in the dataset to avoid an imbalance between classes [25] leading to suboptimal performance.
@@ 3.6 Training and Evaluation Setup
All models were trained on NVIDIA Tesla T4 GPUs with a driver version of 535.104.05 and CUDA version 12.2. All models used a batch size of 4, and were trained using the Stochastic Gradient Descent (SGD) optimizer with a learning rate of 0.001, Nesterov momentum of 0.9, and weight decay of 1e-4.
The models were evaluated using standard objection detection metrics, namely mean Average Precision (mAP) at 0.5 IoU threshold, and Mean Precision [28].
An early stopping mechanism were used to allow each model variant to train for as long as it can improve the validation metrics. At the end of each training epoch, the latest model checkpoint were evaluated against the validation set using the mean Average Precision (mAP) metric. Whenever an improvement in mAP was detected, which corresponds to better model accuracy, the model’s checkpoint was saved as the new best configuration. The training process was stopped if the metrics did not improve over a consecutive span of 10 epochs. However, the checkpoint with the best evaluation metrics were always saved and used for a final evaluation over the test set, which produced the final mAP and Mean Precision metrics for that variant. This adaptive strategy helps avoid overfitting to the training set.
@@
Рис. 5: Comparison of the Depth-only, RGB-only, and RGB-D model variants over mAP and Mean Precision metrics on the test set. The error bars show standard deviation of results over 10 runs.
@@ 4 Results
Рис. 6: Qualitative comparison of the three model variants over two sample scenes from the test set. The RGB-only model tends to fail to detect metallic objects whose color hues are close to the background. The Depth-only model tends to fail detecting thin objects which may not have enough points captured in the 3D point cloud, and it may hallucinate detecting non-existing objects. The RGB-D model shows superior detection accuracy in these examples.
@@
Figure 6 shows the performance of three model variants on two example scenes from the test set arranged side by side for a direct comparison. The qualitative results reinforce the quantitative findings. The RGB-D detections show a notable improvement in identifying objects that the RGB-only model misses–typically metallic objects whose colors are not very different from the scene background. Objects that are misdetected or completely undetected in the RGB-only model are accurately identified by the RGB-D model. This visual comparison aligns with the quantitative results and emphasizes how the integration of depth information with RGB data enhances detection accuracy. The depth component aids in overcoming the limitations observed with RGB-only detections, particularly in complex scenarios where depth cues are crucial for accurate object localization and classification. The Depth-only model’s performance is notably weaker compared to the RGB-only and RGB-D models. The Depth-only model tends to fail detecting objects which are short in height and small objects which may not be captured by many points in the point cloud from the 3D sensor. While depth information can be crucial for adding spatial context in a multimodal setting, it is not sufficient to reliably identify and localize objects, particularly in complex scenes or where depth data lacks the necessary detail and contrast to differentiate objects from the background.
@@ 5 Conclusions
This work demonstrates a notable advancement in the field of object detection for manufacturing by presenting an enhanced multimodal approach that combines RGB and depth data in an efficient manner. This work developed and validated an object detection model using an early fusion of depth information with standard RGB data that results in an efficient four-channel input that can be processed using standard convolutional networks. The experiments have demonstrated that the integration of depth information with RGB data can significantly improve object detection performance and advance the capabilities of object detection systems, particularly in complex manufacturing environments where visual ambiguity and presence of varied objects and components can pose problems for image-only object detection systems. Future work will integrate and evaluate this object detection framework in a robotic object manipulation task [29]. Alternative designs for multimodal object detection will also be explored and evaluated, including the augmentation of 3D detection models with RGB data from cameras.
@@
  • [30] Davis, Jim, Edgar, Thomas, Graybill, Robert, Korambath, Prakashan, Schott, Brian, Swink, Denise, Wang, Jianwu and Wetzel, Jim. ‘‘Smart manufacturing.’’ Annual review of chemical and biomolecular engineering Vol. 6 (2015): pp. 141–160.
  • [31] Mittal, Sameer, Khan, Muztoba Ahmad, Romero, David and Wuest, Thorsten. ‘‘Smart manufacturing: Characteristics, technologies and enabling factors.’’ Proceedings of the Institution of Mechanical Engineers, Part B: Journal of Engineering Manufacture Vol. 233 No. 5 (2019): pp. 1342–1361.
  • [32] Sun, Chen, Shrivastava, Abhinav, Singh, Saurabh and Gupta, Abhinav. ‘‘Revisiting unreasonable effectiveness of data in deep learning era.’’ Proceedings of the IEEE international conference on computer vision: pp. 843–852. 2017.
  • [33] Zhuang, Fuzhen, Qi, Zhiyuan, Duan, Keyu, Xi, Dongbo, Zhu, Yongchun, Zhu, Hengshu, Xiong, Hui and He, Qing. ‘‘A comprehensive survey on transfer learning.’’ Proceedings of the IEEE Vol. 109 No. 1 (2020): pp. 43–76.
  • [34] Sola, J. and Sevilla, J. ‘‘Importance of input data normalization for the application of neural networks to complex industrial problems.’’ IEEE Transactions on Nuclear Science Vol. 44 No. 3 (1997): pp. 1464–1468. 10.1109/23.589532.

### v2

@@ 1 Introduction
Employing Artificial Intelligence (AI) for automation of manufacturing has resulted in increased efficiency, precision, and flexibility and created a paradigm shift in the design of manufacturing systems. AI has been successfully applied to a vast array of manufacturing tasks in the industry [1, 2, 3]. While AI-based methods have improved the manufacturing process, there are still challenges in ensuring that the AI-based black box systems continue to be reliable and robust. The most fundamental layer underlying all smart manufacturing systems is object detection, which allows the system to identify the type and position of the objects that it needs to handle. Object detection is a established computer vision problem, which involves identifying and categorizing specific objects of interest within a larger image by placing a bounding box around each detected object [4].
An effective automation system also requires proper sensor design to provide adequate coverage over the environment and allow the system to properly observe the scene and the objects [5, 6]. The past decades have seen great advancement in sensor hardware. Camera resolutions have increased and they have become more affordable at the same time. Similarly, consumer applications have facilitated mass manufacturing of 3D sensors like lidars and stereo cameras, which are great sensors for smart manufacturing. Despite these improvements, sensors have inherent limitations rooted in their physics [7]. For example, a single image captured by camera does not carry depth information. 3D sensors capture point cloud data which addresses this issue, but these sensors are typically low resolution and do not provide color information.
An effective and reliable automation system requires selecting the right sensors and an object detection system that can effectively ingest the data provided by these sensors. Prior work has shown the limitations of object detection systems that rely solely on cameras [8]. Image distortions like blur and noise can significantly lower the detection accuracy [9]. Although cameras provide color information, there are environments where there is low contrast between objects and the background, and as a result detection accuracy may suffer [10] and cameras may not be enough for handling these environments. Moreover, camera-based object detection systems are sensitive to illumination and can be fragile if there are changes to the lighting conditions. It has also been shown that illumination can negatively affect [11] camera-based object detection systems, because they struggle to generalize to operate under lighting conditions different from what they have experienced during training. Similarly, the performance of these systems diminishes in scenarios where objects vary significantly in size or when they blend indistinguishable with the background in camera’s view [12].
@@
However, none of the existing methods have explored efficient four-channel RGB+D inputs in the context of object detection tasks[25]. The processing of RGB and depth data through a single shared backbone offers several advantages. Firstly, when RGB and depth data are stacked at the input level, the model is able to efficiently extract features that merge information from both modalities, for example color and texture from RGB alongside spatial and structural details from depth. This integrated approach enables the network to learn more meaningful features early in the feature extraction process. By processing both data types simultaneously, the network can better generalize to unseen data as it learns to recognize and interpret patterns across both modalities. Additionally, employing a single backbone for processing reduces the computational overhead significantly. This computational efficiency comes from managing only one set of weights during back-propagation and inference. This approach creates a leaner model which is faster and less costly to run in a manufacturing plant. Lastly, the lower architectural complexity and lack of complex fusion mechanisms simplifies the development and tuning process for training and deploying the final model in the manufacturing environment.
This work introduces a novel method that employs early sensor fusion, where data from RGB images and 3D point clouds are integrated at the data level before being used by the detection algorithm. This method utilizes the combined strengths of both modalities for a more comprehensive representation of the scene that enhances object detection capabilities. Early fusion facilitates the extraction of comprehensive features that embody both the visual and spatial attributes of objects, which addresses the challenges posed by occlusions, variable sizes, and complex background information more effectively than late fusion or single-sensor methods. Hence, this work presents a novel model tailored for processing RGB-D images for object detection tasks in manufacturing. The rest of the paper is organized as follows. Section  2 discusses the methodology of the multimodal object detection framework. Section  3 outlines experiments which compare this approach to both RGB-only and Depth-only object detection of a modified NIST manufacturing task board. Section  4 discusses the experiment results, and Section  5 discusses the conclusions and suggestions for future work.
@@ 3.1 Assembly Task Board
The experiments were designed within the context of robotic pick-and-place applications, which is a common task in industrial manufacturing. The presented framework is relevant as it enables precise localization of the objects, which is critical for grasping and placement in challenging scenarios. The task board used in this work is derived from the modified NIST manufacturing task board configuration which contains nine components serving as the training and testing ground for the proposed object detection system. This board which is shown in Figure 2 includes a large gear, a small gear, a USB-C connector, a nut, a waterproof connector, pairs of small and large rectangular pins, and pairs of small and large round pins. The small and large round pins have diameters of approximately 0.47"and 0.62 with corresponding exposed heights of 1.17"and 1.19 while the rectangular pins measure $0.39"\times 0.62"\times 1.19"$ for the large variant and $0.31"\times 0.47"\times 1.14"$ for the small variant. The gears also differ in size, with the large gear having a diameter of 2.42"and the small gear 1.64 with corresponding exposed heights of 0.69"and 0.66". The USB-C and waterproof connectors have irregular shapes, with bounding box dimensions of $0.25"\times 0.47"\times 1.45"$ for the USB-C connector and $1.27"\times 2.51"\times 0.99"$ for the waterproof connector. The hexagonal nut has a flat-to-flat width of 0.93 an inner diameter of 0.54 and exposed height of 0.43". The usage of this task board introduces realistic scenarios where lighting conditions, component orientation, and material finishes can affect the accuracy of object identification. While the experiments did not include varying lighting conditions, lighting effects were accounted for by incorporating shiny objects such as the four pins and the nut into the dataset. This selection of components provides a comprehensive framework for evaluating the capabilities of the system under study. Variations in surface texture, from gear smoothness to the metallic finish of connectors and pins, generate reflections and shadows that may obscure features or mimic other objects, risking misclassification. In addition, material properties including pin translucency and intricate details of components such as USB-C connectors which have very few points in space that represents its depth introduce visual variability that demands sophisticated interpretation by the network. These factors underscore the need for an advanced, adaptable neural network and thorough pre-processing to ensure accurate detection amidst the multifaceted visual conditions of the task board.
@@ 3.2 Sensors and Calibration
The object detection framework ingests data from two independent sensors capturing image and depth data. Figure 3 shows the multi-sensor setup used in our experiments.The sensors are mounted together and affixed to the robot’s table. A Basler camera captures high resolution images to provide detailed texture, color, and shape information which are necessary for distinguishing objects based on color contrast against the background. An Intel RealSense depth camera senses the environment 3D in the form of point clouds to capture the spatial structure of the scene. This additional depth information allows the model to better differentiate objects and estimate their positions. This combination of sensors was selected to provide visual details of the scene in addition to the spatial relationships and dimensions of the objects under study. The separate Basler RGB camera was used to ensure higher image quality and flexibility in the multimodal setup. It has higher resolution and color accuracy in comparison to the RealSense RGB sensor. Although certain types of hardware such as RealSense used in this study can provide both RGB and depth data from a single sensor, our multi-sensor framework provides greater compatibility and generalizability to diverse manufacturing problems and environments. Using this approach any kind of sensor with any combinations can be integrated to satisfy needs of the manufacturing environment.
The two sensors need to be calibrated against each other since the assembly task and the object detection model require precise alignment between the image and depth data received from the environment. Camera calibration [27] involves the precise estimation of camera parameters, including intrinsic and extrinsic, to infer accurate geometric features from captured sequences. To conduct the calibration, an asymmetric 10x7 checkerboard pattern was placed in various positions and orientations within the two sensor’s field of view and 60 pairs of image and point clouds were captured. MATLAB’s Lidar-Camera Calibration application [28] was used to tune the calibration parameters, as shown in Fig. 4. Fig. 4 shows that the camera intrinsics are properly calibrated, therefore the red boundary perfectly lines up with the checkerboard’s perimeter. However, the extrinsics are not calibrated yet, so the point cloud does not align with the checkerboard in the camera view. At the end of the process, the visual display and the error metrics signal a proper calibration. Hence, this process involved iteratively removing image and point cloud pairs with high calibration errors until the maximum translation error was below $0.0045m$ and the maximum rotation error was below $4.5\degree$. At this point, a total of 37 image - point cloud pairs were used to compute the calibration parameters.
Рис. 3: Multi-sensor setup featuring a Basler RGB camera and an Intel RealSense camera mounted to the robot’s table. The location and orientation of the sensors remained fixed throughout all experiments.
Рис. 4: MATLAB’s Lidar-Camera Calibration application output before 3D point cloud alignment.
@@ 3.3 Generating Depth Maps
Since the model receives RGB images encoded as unsigned integers in range $[0,255]$, this work further scales the depth values as
@@
where $D_{s}$ is the scaled depth value. After this transformation, the depth map is concatenated with the raw RGB image to form a four-channel image where all values range consistently from 0 to 255, which makes the data suitable for feeding to the model. The four-channel RGB-D images can be conveniently precomputed and stored to disk since many image formats like PNG support a fourth alpha channel mainly used for encoding transparency information for each pixel.
@@ 3.4 Data Collection and Labeling
A dataset containing 301 pairs of images each in 1920 pixels in width and 1200 pixels in height, along with associated point clouds captured from varied configurations of the assembly components on and off the assembly board were created. Placing the objects on and off the board enables the robot to detect and localize the components both before and after they get installed on the assembly board. The position and orientation of the assembly board within the sensors’ field of view were also varied so that the model can generalize to varied task configurations. It was also ensured that each class of components is adequately present in the dataset to avoid an imbalance between classes [29] leading to suboptimal performance.
@@ 3.5 Model Variants
Рис. 5: The three variants of the object detection model used in experiments. a. RGB-only: The model receives only camera images and predicts object class scores and bounding boxes for the detected objects. b. RGBD-Man: The main model which concatenates RGB and depth information into a four-channel input and predicts objects from the fused inputs. c. Depth-only: The model which predicts objects given only a single-channel depth map of scene.
@@ 3.6 Training and Evaluation Setup
All models were trained on NVIDIA Tesla T4 GPUs with a driver version of 535.104.05 and CUDA version 12.2. All models used a batch size of 4, which was a compromise between ensuring sufficient gradient stability and keeping the memory requirements within the limits of our hardware. The models were trained using the Stochastic Gradient Descent (SGD) optimizer with a learning rate of 0.001, Nesterov momentum of 0.9, and weight decay of 1e-4. SGD has been widely used in object detection tasks and has proven effective in training Faster R-CNN models. The SGD optimizer with the specified learning rate and Nesterov momentum helps the model converge to an appropriate local minimum while also benefiting from faster updates due to the momentum term. The SGD optimizer is computationally efficient, and can lead to better generalization performance compared to other optimizers like Adam.
The models were evaluated using standard objection detection metrics, namely mean Average Precision (mAP) at 0.5 IoU threshold, and Mean Precision [32]. The 0.5 IoU threshold for mAP is a widely accepted metric for evaluating the model performance in object detection benchmarks. This threshold provides a balance between localization accuracy and detection tolerance. It ensures that detected bounding boxes reasonably overlap with ground truth annotations.
An early stopping mechanism was used to allow each model variant to train for as long as it can improve the validation metrics. At the end of each training epoch, the latest model checkpoint were evaluated against the validation set using the mean Average Precision (mAP) metric. Whenever an improvement in mAP was detected, which corresponds to better model accuracy, the model’s checkpoint was saved as the new best configuration. The training process was stopped if the metrics did not improve over a consecutive span of 10 epochs. However, the checkpoint with the best evaluation metrics was always saved and used for a final evaluation over the test set, which produced the final mAP and Mean Precision metrics for that variant. This adaptive strategy helps avoid overfitting to the training set, and validates the chosen hyperparameters, including batch size and optimizer. The consistency of the results across multiple runs further supports the reliability of the training setup.
@@
Рис. 6: Comparison of the Depth-only, RGB-only, and RGB-D model variants over mAP and Mean Precision metrics on the test set. The error bars show standard deviation of results over 10 runs.
@@ 4 Results
Рис. 7: Qualitative comparison of the three model variants over two sample scenes from the test set. The RGB-only model tends to fail to detect metallic objects whose color hues are close to the background. The Depth-only model tends to fail detecting thin objects which may not have enough points captured in the 3D point cloud, and it may hallucinate detecting non-existing objects. The RGB-D model shows superior detection accuracy in these examples.
@@
Figure 7 shows the performance of three model variants on two example scenes from the test set arranged side by side for a direct comparison. The qualitative results reinforce the quantitative findings. The RGB-D detections show a notable improvement in identifying objects that the RGB-only model misses–typically metallic objects whose colors are not very different from the scene background. The RGB-only model also struggles with reflective surface objects and objects with low contrast with the background. Metal pins often are missed in detection because reflection causes inconsistent pixel values. These corresponding RGB values change with different orientations and lighting, which lead to misdetections, as the model cannot reliably differentiate object boundaries. Objects that are misdetected or completely undetected in the RGB-only model are accurately identified by the RGB-D model. This visual comparison aligns with the quantitative results and emphasizes how the integration of depth information with RGB data enhances detection accuracy. The depth component aids in overcoming the limitations observed with RGB-only detections, particularly in complex scenarios where depth cues are crucial for accurate object localization and classification. The Depth-only model’s performance is notably weaker compared to the RGB-only and RGB-D models. The Depth-only model tends to fail detecting objects which are short in height and small objects which may not be captured by many points in the point cloud from the 3D sensor. Note that the model lacks a full 3D representation or point cloud data. Only the depth values for every pixel in each image was given to the model in the shape of a one dimensional array. This results in the model lacking texture, shape, and edge information that are critical for object recognition. While depth information can be crucial for adding spatial context in a multimodal setting, it is not sufficient to reliably identify and localize objects, particularly in complex scenes or where depth data lacks the necessary detail and contrast to differentiate objects from the background.
@@ 5 Conclusions
This work demonstrates a notable advancement in the field of object detection for manufacturing by presenting an enhanced multimodal approach that combines RGB and depth data in an efficient manner. This work developed and validated an object detection model using an early fusion of depth information with standard RGB data that results in an efficient four-channel input that can be processed using standard convolutional networks. The experiments have demonstrated that the integration of depth information with RGB data can significantly improve object detection performance and advance the capabilities of object detection systems, particularly in complex manufacturing environments where visual ambiguity and presence of varied objects and components can pose problems for image-only object detection systems.
Although there is still a gap between the current results and the high accuracy required for industrial application, the performance would likely improve by utilizing a larger dataset with more diverse object configurations, lighting conditions, and occlusion scenarios. Additionally, a better sensor calibration and refined depth processing alongside with transfer learning and fine tuning the model on industrial settings will also help improve the accuracy to meet real world manufacturing requirements. Future work will explore scenarios involving transparent, irregular, and non-standard shapes, as well as fully 3D and overlapping objects to further investigate the role of depth in enhancing object detection in manufacturing environments. Moreover, future work will integrate and evaluate this object detection framework in a robotic object manipulation task [33]. As future work, controlled robotic grasping experiments will be performed to determine if the current detection performance is sufficient for this robotic grasping task, or whether further refinements are necessary. Alternative designs for multimodal object detection will also be explored and evaluated, including the augmentation of 3D detection models with RGB data from cameras.
@@
  • [3] Ziad, Erfan, Yang, Zhuo, Lu, Yan and Ju, Feng. ‘‘Knowledge Constrained Deep Clustering for Melt Pool Anomaly Detection in Laser Powder Bed Fusion.’’ 2024 IEEE 20th International Conference on Automation Science and Engineering (CASE): pp. 670–675. 2024. IEEE.
@@
  • [6] Ziad, Erfan, Ju, Feng, Yang, Zhuo and Lu, Yan. ‘‘Pyramid Learning Based Part-to-Part Consistency Analysis in Laser Powder Bed Fusion.’’ International Manufacturing Science and Engineering Conference, Vol. 88100: p. V001T01A024. 2024. American Society of Mechanical Engineers.
  • [7] Jang, Youjin, Jeong, Inbae, Younesi Heravi, Moein, Sarkar, Sajib, Shin, Hyunkyu and Ahn, Yonghan. ‘‘Multi-Camera-Based Human Activity Recognition for Human–Robot Collaboration in Construction.’’ Sensors Vol. 23 No. 15 (2023). 10.3390/s23156997. URL https://www.mdpi.com/1424-8220/23/15/6997.
@@
  • [25] Zhou, Tao, Fan, Deng-Ping, Cheng, Ming-Ming, Shen, Jianbing and Shao, Ling. ‘‘RGB-D salient object detection: A survey.’’ Computational Visual Media Vol. 7 (2021): pp. 37–69.